OpenStudy (anonymous):

The second hand of a clock sweeps continuously around the face of the clock. What is the probability that at any random moment the second hand is between 7 & 12? Pleaseeeee Help :)

OpenStudy (anonymous):

Strictly between 7 and 12, given some random interval in time? Then: \[ 60-(12-7)5-1=35-1=34 \]We have 60 seconds in total, so: \[ P(E)=\frac{34}{60}=\frac{17}{30} \]And that is your probability.

OpenStudy (anonymous):

Thanks :))

OpenStudy (anonymous):

Sure thing

OpenStudy (unklerhaukus):

5/12

OpenStudy (anonymous):

depends on what time it is.

OpenStudy (anonymous):

@UnkleRhaukus how'd you'd get 5/12? cuz thats the right answer

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

wolf

OpenStudy (anonymous):

My God I'm stupid, jeez, I literally just gave you the inverse of the proability. It'd be \[ \frac{5\cdot5-1}{60}=\frac{24}{60}=\frac{2}{5} \]If it is strictly between, or, if it includes 7, then\[\frac{12-7}{12}=\frac{5}{12}\]

OpenStudy (anonymous):

I need to get some sleep...

OpenStudy (anonymous):

lol thanks @LolWolf

OpenStudy (unklerhaukus):

the probability that the hand in on the clock face is 12/12 the probability the hand will be between adjacent numbers is 1/12 the probability the hand will be within 5 adjacent numbers is 5/12